OpenStudy (anonymous):

A sled is being pulled across a barn floor by a horizontal force of +320 N. The coefficient of sliding friction between the sled and the florr is 0.45. The mass of the sled is 65 kg. Determine the Normal Force that the barn floor exerts on the sled as well as the acceleration of thhe sled..... Please Help!?!

OpenStudy (fretje):

assuming the floor being horizontal, the weight is the normal force G = m.g = 65kg.9,8m/s^2 = 637N [kg.m/s^2] according to the second law of motion of newton. for sliding object horizontal force is \[F _{horiz} = \phi _{dyn}. G\] here phi represents dynamic friction coefficient (static friction coefficient can be different but is here irrelevant since the sled is allready moving) so i get Fhoriz = 0,45 . 637N = 286.65 N I have a net Force of F = Fpull-Fhoriz = 33.35 N this force is used to accelerate the sled of 65 kg again second law F = m.a a = F/m = 33,35N/65kg = 0.513m/s^2

OpenStudy (anonymous):

Thank you so much!

OpenStudy (fretje):

i like you anyway

OpenStudy (anonymous):

would you mind helping me with a couple of other questions?

OpenStudy (fretje):

i have little time, but no, i can help

OpenStudy (fretje):

sorry, my friend came, we go home (live in belgium). But give me the questions here, and i answer them tomorrow

OpenStudy (anonymous):

ok... A person weighs 1500 N and is standing in an elevator on a scale that is currently reading 1200 N as the are descending. What is happening to this person? If the elevator is accelerating, determine this acceleration.

OpenStudy (anonymous):

alright that's fine. thanks anyway! (:

OpenStudy (fretje):

So the person reads a force of 1200N on the scale. This is less than his weight on earth, which is 1500N, or thus having a mass of 1500N/9,8m/s²=153kg (mass m = F/g with g the gravitational acceleration 9,8m/s²). The person feels lighter, less pressure in the legs, than on earth, like he/she is falling. This can only mean two things: either he/she accelerates towards a celestial body, with a velocity greater and greater. The second law of motion of Newton dictates: F= m.a. The force F is the net force on his/her body exerted by the scale on him/her 1500N-1200N = 300N The acceleration is therefore a=F/m=300N/153kg = 1,96 m/s² Or second possibility: he/she is not in a gravitational field, and is accelerated upwards with a = 9,8m/s² - 1,96m/s² = 7,84m/s²

OpenStudy (fretje):

but the last possibility I should rule out, as in the question is said that the person is descending.

OpenStudy (anonymous):

thankyou! you're awesome! :)